# Space spanned by characteristic exponents

### v2

@@ 5.2. Linear algebra for multiplication
For every field $k$ of characteristic 00, denote by $\mu_{k}$ the subgroup of roots of unity in $k$. Denote by
@@
the quotient map. Extend ${\rm rog}$ to a map ${\rm rog}:k\to\mathbb{D}(k)\cup\{\infty\}$ by sending 00 to $\infty$. Here we use the notation “${\rm rog}$” since it is an analogy of the classical $\log$ function to some extent. The embedding $k\hookrightarrow\overline{k}$ gives a natural embedding $\mathbb{D}(k)\hookrightarrow\mathbb{D}(\overline{k})$ as multiplicative abelian groups. Set $\mathbb{D}(k)_{\mathbb{Q}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{Q}$, where $\mathbb{D}(k)$ is viewed as a multiplicative abelian group, hence a $\mathbb{Z}$-module. Then $\mathbb{D}(k)_{\mathbb{Q}}$ is the subspace of $\mathbb{D}(\overline{k})$ spanned by $\mathbb{D}(k)$ over $\mathbb{Q}$. Set $\mathbb{D}(k)_{\mathbb{R}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{R}$.
@@ 5.3. Norms
Let $k$ be a field of characteristic 00. For every finite field extension $\widetilde{k}$ of $k$, denote by $N_{\widetilde{k}/k}:\widetilde{k}\to k$ the norm map. The norm map induces a map
$\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{Q}}\to\mathbb{D}(k)_{\mathbb{Q}},{\rm rog}(x)\mapsto[l:k]^{-1}{\rm rog}(N_{l/k}(x)),$
where $l$ is any finite field extension of $k$ containing $x$. One can check that $\mathbf{n}_{k}$ is well defined and is $\mathbb{Q}$-linear. We also denote by $\mathbf{n}_{k}$ its $\mathbb{R}$-linear extension $\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{R}}\to\mathbb{D}(k)_{\mathbb{R}}$. When the field $k$ is clear, we also write $\mathbf{n}$ for $\mathbf{n}_{k}$.
@@ 5.4. Valuations
Assume that $K$ is a number field. Denote by $\mathcal{M}_{K}$ the set of all places of $K$. For every $v\in\mathcal{M}_{K}$, we define a map $\phi_{v}:\mathbb{D}(K)_{\mathbb{R}}\to\mathbb{R}$, by $\phi_{v}({\rm rog}(x))=-\log(\lvert x\rvert_{v})$ for $x\in K^{*}$ and extending $\mathbb{R}$-linearly, where $\left|\cdot\right|_{v}$ is a fixed multiplicative norm on $K$ corresponding to $v$. It is easy to check that the map $\phi_{v}$ is well-defined and $\mathbb{R}$-linear. We denote the restriction of $\phi_{v}$ to $\mathbb{D}(K)_{\mathbb{Q}}$ by $\phi_{v}:\mathbb{D}(K)_{\mathbb{Q}}\to\mathbb{R}$ as well. For every $a\in\mathbb{D}(K)_{\mathbb{R}}$, it is clear that the set $\{v\in\mathcal{M}_{K}:\phi_{v}(a)\neq 0\}$ is finite.
Let $S$ be a finite subset of $\mathcal{M}_{K}$ containing all the archimedean places. Let $\mathcal{O}_{K,S}$ be the ring of $S$-integers in $K$. Let $\mathcal{O}$ be the integral closure of $\mathcal{O}_{K,S}$ in $\overline{K}$. For every $v\in\mathcal{M}_{K}\setminus S$ and $\lambda\in\mathcal{O}$, we have $\phi_{v}\circ\mathbf{n}_{K}(\lambda)\geq 0$. Write $\mathcal{M}_{K}\setminus S=\{v_{1},v_{2},\dots\}$. Then $(\phi_{v_{i}})_{i=1}^{\infty}\subseteq{\rm Hom}(\mathbb{D}(K)_{\mathbb{R}},\mathbb{R})$ is an $\mathcal{O}_{K,S}$-sequence and $(\phi_{v_{i}}\circ\mathbf{n}_{K})_{i=1}^{\infty}\subseteq{\rm Hom}(\mathbb{D}(\overline{K})_{\mathbb{R}},\mathbb{R})$ is an $\mathcal{O}$-sequence.
@@ 5.5. Complex conjugation and absolute value
Let $\mathbf{k}$ be an algebraically closed subfield of $\mathbb{C}$ stable under the complex conjugation $\tau$. Still denote by $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ the restriction of the complex conjugation on $\mathbf{k}$, which is an involution of $\mathbf{k}$. Denote by $\mathbf{k}^{\tau}=\mathbf{k}\cap\mathbb{R}$ the $\tau$-invariant subfield of $\mathbf{k}$. Then the restriction of the absolute value $\mathbf{n}_{\mathbb{R}}$ on $\mathbf{k}$ is
@@
We shall prove the following theorem involving an involution $\tau$ (which may not be the complex conjugation and may be the identity) in the next section.
@@
Let $\mathbf{k}$ be an algebraically closed field of characteristic 00. Let $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ be an element with $\tau^{2}=\mathrm{id}$. If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace in $\mathbb{D}(\mathbf{k}^{\tau})_{\mathbb{Q}}$ spanned by $\{\mathbf{n}_{\mathbf{k}^{\tau}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
Applying Theorem 5.7 to the case where $\mathbf{k}=\mathbb{C}$ and $\tau$ is the complex conjugation, we deduce Theorem 1.4 in the case where $f$ is not PCF.
@@
Setting $\tau=\mathrm{id}$, then from Theorem 5.7 we get the following result:
Assume that $\mathbf{k}$ is an algebraically closed field of characteristic 00. If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace in $\mathbb{D}(\mathbf{k})_{\mathbb{Q}}$ spanned by $\{{\rm rog}(\rho_{f}(x)):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
@@
Denote by $\mathcal{C}_{f}$ the set of critical points of $f$. Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ such that the (forward) $f$-orbit $O_{f}(o)$ of $o$ is infinite. We fix such a critical point $o$. Let $X$ be the union of all (forward) orbits of periodic critical points of $f$. Then $X$ is finite.
@@
Denote by $\mathcal{M}_{K}$ the set of all places of $K$. Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all the archimedean places, satisfying $\tau(B)=B$, such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$. Then we have $\tau(\mathcal{O}_{K,B})=\mathcal{O}_{K,B}$. For $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$, set $\lambda(x)=(n_{f}(x))^{-1}{\rm rog}(\rho_{f}(x))\in\mathbb{D}(\overline{\mathbb{Q}})_{\mathbb{R}}\cup\{\infty\}$. Recall that $n_{f}(x)\in\mathbb{Z}_{>0}$ is the exact period of $x$ and $\rho_{f}(x)=df^{n_{f}(x)}(x)\in\overline{\mathbb{Q}}$ is the multiplier of $x$. For all $x\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$, we have $\mathbf{n}_{K}(\lambda(x))\in O:=\mathbb{D}(\mathcal{O}_{K,B})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ by [24, Corollary 2.23 (a)]. Note that $O=\tau(O)$ is invariant under $\tau$, where $\tau$ descends to $\mathbb{D}(K)$ and is extended $\mathbb{Q}$-linearly to $\mathbb{D}(K)_{\mathbb{Q}}$.
@@
For every $v\in\mathcal{M}_{K}\setminus B$ and $x\in\mathbb{P}^{1}(\mathbb{C}_{v})$, denote by $\tilde{x}\in\mathbb{P}^{1}(\overline{\widetilde{K_{v}}})$ the reduction of $x$ in the special fiber at $v$ and $f_{v}:\mathbb{P}^{1}_{\overline{\widetilde{K_{v}}}}\to\mathbb{P}^{1}_{\overline{\widetilde{K_{v}}}}$ the reduction of $f$. (Here, $\overline{\widetilde{K_{v}}}$ is a fixed algebraic closure of the residue field $\widetilde{K_{v}}$ of $K_{v}$.) After enlarging $B$, we can assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$, where $X_{v}$ is the reduction of $X$ contained in $\mathbb{P}^{1}(\widetilde{K_{v}})\subseteq\mathbb{P}^{1}(\overline{\widetilde{K_{v}}})$.
@@
(iii) there exist an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$, $q\in\mathcal{C}_{f}$, and $m\in\mathbb{Z}_{\geq 0}$, such that $\widetilde{\sigma(q)}$ is $f_{v}$-periodic with $\widetilde{\sigma(x)}=f_{v}^{m}(\widetilde{\sigma(q)})$.
For every $v\in\mathcal{M}_{K}\setminus B$, denote by $P_{v}$ the union of all $f_{v}$-orbits of periodic critical points of $f_{v}$. The set $P_{v}$ is finite. For every $v\in\mathcal{M}_{K}\setminus B$ and $q\in P_{v}$, by Hensel’s lemma there exists a unique periodic point $y=y(q)\in{\rm Per}(f)(\mathbb{C}_{v}\cap\overline{\mathbb{Q}})$ such that $\tilde{y}=q$. Then there exists a unique ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit $O(q)$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}})$ such that for some (then every) $x\in O(q)$, there exists an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ satisfying $\widetilde{\sigma(x)}=q$ (here $O(q)$ is the ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit of $y=y(q)$). In particular, we have $X_{v}\subseteq P_{v}$ and $\cup_{q\in X_{v}}O(q)=X$. It follows that the set
@@
Recall that $o\in\mathcal{C}_{f}$ is a critical point of $f$ which is not $f$-preperiodic and we assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$. By [3, Lemma 4.1], there are infinitely may $v\in\mathcal{M}_{K}\setminus B$, such that there exists $n\in\mathbb{Z}_{>0}$ with $f_{v}^{n}(\tilde{o})=\tilde{o}$. For such a place $v$, we have $\tilde{o}\in P_{v}\setminus X_{v}$, which concludes the proof. ∎
@@
There exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{i}}(\mathbf{n}(\lambda(x_{i})))>0$ for all $i\geq 1$ and $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))=0$ for $j\neq i$. In particular, $((\mathbf{n}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangle $O$-system for $\mathbb{D}(K)_{\mathbb{Q}}$.
@@
We construct these two sequences recursively.
By Lemma 6.1, $S$ is infinite. Pick $v_{1}\in S$, then there exists $x_{1}\in Q_{v_{1}}\setminus X\subseteq$ ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$. We have $\phi_{v_{1}}(\mathbf{n}(\lambda(x_{1})))>0$.
Assume that we have constructs $x_{1},\dots,x_{m}\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and $v_{1},\dots,v_{m}\in$ $\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))\geq 0$ and the equality holds if and only if $j\neq i$. The set $\cup_{i=1}^{m}Q_{v_{i}}\setminus X$ is finite. Then there exists a finite set $T_{m}\subseteq\mathcal{M}_{K}$ such that for all $x\in\cup_{i=1}^{m}Q_{v_{i}}\setminus X$ and $v\in\mathcal{M}_{K}\setminus T_{m}$, we have $\phi_{v}(\mathbf{n}(\lambda(x)))=0$. By Lemma 6.1, we can take a place $v_{m+1}\in$ $S\setminus(\{v_{1},\dots,v_{m}\}\cup T_{m})$. Then we have $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{i})))=0$ for all $1\leq i\leq m$. Pick $x_{m+1}\in Q_{v_{m+1}}\setminus X$. We have $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{m+1})))>0$. It follows that $x_{m+1}\notin\cup_{i=1}^{m}Q_{v_{i}}$. Thus $\phi_{v_{i}}(\mathbf{n}(\lambda(x_{m+1})))=0$ for all $1\leq i\leq m$. We conclude the proof of Lemma 6.2. ∎
Then we conclude the proof by Corollary 5.5 and Remark 5.6.∎
@@
Denote by $\mathcal{C}_{f}$ the set of critical points of $f$. Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ which is not $f$-preperiodic. We fix such a critical point $o$. Fix a subfield $K$ of $\mathbf{k}$ such that $K/\mathbb{Q}$ is finite generated, $\tau(K)=K$, and $o,f$ are defined over $K$. Without loss of generality, we may assume that $\mathbf{k}=\overline{K}$.
Take a finite generated $\mathbb{Z}$-subalgebra $A$ of $K$ with ${\rm Frac}(A)=K$ and $\tau(A)=A$. After shrinking ${\rm Spec}(A)$, we may assume that there exists an endomorphism $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ over $A$ whose restriction $f_{K}:\mathbb{P}^{1}_{K}\to\mathbb{P}^{1}_{K}$ on the generic fiber $\mathbb{P}^{1}_{K}$ satisfies $f=f_{K}\otimes_{K}\mathbf{k}$.
For every $c\in{\rm Spec}(A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}})(\overline{\mathbb{Q}})$, denote by $f_{c}$ the specialization of $f_{A}$ at $c$, and $o_{c}$ the specialization of $o$ at $c$. Then $o_{c}$ is a critical point of $f_{c}$. By [9, Lemma 3.3], we can take $c\in{\rm Spec}(A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}})(\overline{\mathbb{Q}})$ such that the $f_{c}$-orbit of $o_{c}$ is infinite, i.e., $o_{c}$ is not $f_{c}$-preperiodic. In particular, the specialization map $f_{c}$ is not PCF as well. We take a number field $L\subseteq\overline{K}$ such that $c$ is defined over $L$ and $\tau(L)=L$. Denote by $A_{1}$ the $\mathbb{Z}$-algebra generated by $A$, $\mathcal{O}_{L}$, and $\tau(\mathcal{O}_{L})$ in $A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}}$. We may replace ${\rm Spec}(A)$ by a suitable non-empty affine Zariski open set of ${\rm Spec}(A_{1})$ such that $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ is still everywhere well-defined and $c\in{\rm Spec}(A\otimes_{\mathcal{O}_{L}}L)(L)$, after the replacement. We view ${\rm Spec}(A)$ as an $\mathcal{O}_{L}$-scheme. After shrinking ${\rm Spec}(A)$, the Zariski closure of $c$ in ${\rm Spec}(A)$ is isomorphic to ${\rm Spec}(\mathcal{O}_{L,S})$ for a finite set $S\subseteq\mathcal{M}_{L}$ of places of $L$ containing all archimedean places, which corresponds to a prime ideal $p$ of $A$.

### v3

@@ 5.2. Linear algebra for multiplication
For every field $k$ of characteristic 0, denote by $\mu_{k}$ the subgroup of roots of unity in $k$. Let
@@
be the quotient map. Extend $\rm rog$} to a map ${\rm rog}:k\to\mathbb{D}(k)\cup\{\infty\}$ by sending 0 to $\infty$. (The notation $\rm rog$}” is used by analogy with the classical $\log$ function.) The embedding $k\hookrightarrow\overline{k}$ induces a natural embedding $\mathbb{D}(k)\hookrightarrow\mathbb{D}(\overline{k})$ of multiplicative abelian groups. Set $\mathbb{D}(k)_{\mathbb{Q}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{Q}$, where $\mathbb{D}(k)$ is viewed as a multiplicative abelian group (hence a $\mathbb{Z}$-module). Then $\mathbb{D}(k)_{\mathbb{Q}}$ is the subspace of $\mathbb{D}(\overline{k})$ spanned by $\mathbb{D}(k)$ over $\mathbb{Q}$. Set $\mathbb{D}(k)_{\mathbb{R}}:=\mathbb{D}(k)\otimes_{\mathbb{Z}}\mathbb{R}$.
@@ 5.3. Norms
Let $k$ be a field of characteristic 0. For every finite extension $\widetilde{k}$ of $k$, denote by $N_{\widetilde{k}/k}:\widetilde{k}\to k$ the norm map. The norm map induces a map
$\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{Q}}\to\mathbb{D}(k)_{\mathbb{Q}},{\rm rog}(x)\mapsto[k^{\prime}:k]^{-1}{\rm rog}(N_{k^{\prime}/k}(x)),$
where $k^{\prime}$ is any finite extension of $k$ containing $x$. One can check that $\mathbf{n}_{k}$ is well-defined and $\mathbb{Q}$-linear. We also denote by $\mathbf{n}_{k}$ its $\mathbb{R}$-linear extension $\mathbf{n}_{k}:\mathbb{D}(\overline{k})_{\mathbb{R}}\to\mathbb{D}(k)_{\mathbb{R}}$. When the field $k$ is clear, we write $\mathbf{n}$ for $\mathbf{n}_{k}$.
@@ 5.4. Valuations
Assume that $K$ is a number field. Denote by $\mathcal{M}_{K}$ the set of all places of $K$. For every $v\in\mathcal{M}_{K}$, we define a map $\phi_{v}:\mathbb{D}(K)_{\mathbb{R}}\to\mathbb{R}$ by $\phi_{v}({\rm rog}(x))=-\log(\lvert x\rvert_{v})$ for $x\in K^{*}$ and extending $\mathbb{R}$-linearly, where $\left|\cdot\right|_{v}$ is a fixed multiplicative norm on $K$ corresponding to $v$. It is easy to check that the map $\phi_{v}$ is well-defined and $\mathbb{R}$-linear. We denote the restriction of $\phi_{v}$ to $\mathbb{D}(K)_{\mathbb{Q}}$ by $\phi_{v}:\mathbb{D}(K)_{\mathbb{Q}}\to\mathbb{R}$ as well. For every $a\in\mathbb{D}(K)_{\mathbb{R}}$, it is clear that the set $\{v\in\mathcal{M}_{K}:\phi_{v}(a)\neq 0\}$ is finite.
Let $S$ be a finite subset of $\mathcal{M}_{K}$ containing all archimedean places. Let $\mathcal{O}_{K,S}$ be the ring of $S$-integers in $K$. Let $\mathcal{O}$ be the integral closure of $\mathcal{O}_{K,S}$ in $\overline{K}$. For every $v\in\mathcal{M}_{K}\setminus S$ and $\lambda\in\mathcal{O}$, we have $\phi_{v}\circ\mathbf{n}_{K}(\lambda)\geq 0$. Write $\mathcal{M}_{K}\setminus S=\{v_{1},v_{2},\dots\}$. Then $(\phi_{v_{i}})_{i=1}^{\infty}\subseteq{\rm Hom}(\mathbb{D}(K)_{\mathbb{R}},\mathbb{R})$ is an $\mathcal{O}_{K,S}$-sequence and $(\phi_{v_{i}}\circ\mathbf{n}_{K})_{i=1}^{\infty}\subseteq{\rm Hom}(\mathbb{D}(\overline{K})_{\mathbb{R}},\mathbb{R})$ is an $\mathcal{O}$-sequence.
@@ 5.5. Complex conjugation and absolute value
Let $\mathbf{k}$ be an algebraically closed subfield of $\mathbb{C}$ stable under complex conjugation $\tau$. Still denote by $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ the restriction of complex conjugation on $\mathbf{k}$, which is an involution of $\mathbf{k}$. Let $\mathbf{k}^{\tau}=\mathbf{k}\cap\mathbb{R}$ be the $\tau$-invariant subfield of $\mathbf{k}$. Then the restriction of the absolute value $\mathbf{n}_{\mathbb{R}}$ on $\mathbf{k}$ is
@@
We shall prove the following theorem involving an involution $\tau$ (which may not be complex conjugation and may be the identity) in the next section.
@@
Let $\mathbf{k}$ be an algebraically closed field of characteristic 0. Let $\tau\in{\rm Gal}(\mathbf{k}/\mathbb{Q})$ be an element with $\tau^{2}=\mathrm{Id}$. If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbf{k}^{\tau})_{\mathbb{Q}}$ spanned by $\{\mathbf{n}_{\mathbf{k}^{\tau}}({\rm rog}(\rho_{f}(x))):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
Applying Theorem 5.7 to the case $\mathbf{k}=\mathbb{C}$ and $\tau$ being complex conjugation, we deduce Theorem 1.4 in the case where $f$ is not PCF.
@@
By setting $\tau=\mathrm{Id}$, Theorem 5.7 yields the following result:
Assume that $\mathbf{k}$ is an algebraically closed field of characteristic 0. If $f:\mathbb{P}^{1}_{\mathbf{k}}\to\mathbb{P}^{1}_{\mathbf{k}}$ is an endomorphism over $\mathbf{k}$ of degree at least $2$ which is not PCF, then the $\mathbb{Q}$-subspace of $\mathbb{D}(\mathbf{k})_{\mathbb{Q}}$ spanned by $\{{\rm rog}(\rho_{f}(x)):x\in{\rm Per}^{*}(f)(\mathbf{k})\}$ is of infinite dimension.
@@
Denote by $\mathcal{C}_{f}$ the set of critical points of $f$. Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ such that the (forward) $f$-orbit $O_{f}(o)$ of $o$ is infinite. Fix such a critical point $o$. Let $X$ be the union of all (forward) orbits of periodic critical points of $f$. Then $X$ is finite.
@@
Denote by $\mathcal{M}_{K}$ the set of all places of $K$. Take a finite subset $B\subseteq\mathcal{M}_{K}$ containing all archimedean places, satisfying $\tau(B)=B$, and such that $f$ has good reduction at $v$ for every $v\in\mathcal{M}_{K}\setminus B$. Then $\tau(\mathcal{O}_{K,B})=\mathcal{O}_{K,B}$. For $x\in{\rm Per}(f)(\overline{\mathbb{Q}})$, set $\lambda(x)=(n_{f}(x))^{-1}{\rm rog}(\rho_{f}(x))\in\mathbb{D}(\overline{\mathbb{Q}})_{\mathbb{R}}\cup\{\infty\}$. Recall that $n_{f}(x)\in\mathbb{Z}_{>0}$ is the exact period of $x$ and $\rho_{f}(x)=df^{n_{f}(x)}(x)\in\overline{\mathbb{Q}}$ is the multiplier of $x$. For all $x\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$, we have $\mathbf{n}_{K}(\lambda(x))\in O:=\mathbb{D}(\mathcal{O}_{K,B})\otimes_{\mathbb{Z}}\mathbb{Q}_{>0}$ by [24, Corollary 2.23 (a)]. Note that $O=\tau(O)$ is invariant under $\tau$, where $\tau$ descends to $\mathbb{D}(K)$ and is extended $\mathbb{Q}$-linearly to $\mathbb{D}(K)_{\mathbb{Q}}$.
@@
[…]
@@
[…]
For every $v\in\mathcal{M}_{K}\setminus B$, denote by $P_{v}$ the union of all $f_{v}$-orbits of periodic critical points of $f_{v}$. The set $P_{v}$ is finite. For every $v\in\mathcal{M}_{K}\setminus B$ and $q\in P_{v}$, by Hensel’s lemma there exists a unique periodic point $y=y(q)\in{\rm Per}(f)(\mathbb{C}_{v}\cap\overline{\mathbb{Q}})$ such that $\tilde{y}=q$. Then there exists a unique ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit $O(q)$ in $\mathbb{P}^{1}(\overline{\mathbb{Q}})$ such that for some (hence every) $x\in O(q)$, there exists an embedding $\sigma:\overline{\mathbb{Q}}\hookrightarrow\mathbb{C}_{v}$ satisfying $\widetilde{\sigma(x)}=q$ (here $O(q)$ is the ${\rm Gal}(\overline{\mathbb{Q}}/K)$-orbit of $y=y(q)$). In particular, we have $X_{v}\subseteq P_{v}$ and $\cup_{q\in X_{v}}O(q)=X$. It follows that the set
@@
Recall that $o\in\mathcal{C}_{f}$ is a critical point of $f$ that is not $f$-preperiodic, and we assume that $\tilde{o}\notin X_{v}$ for every $v\in\mathcal{M}_{K}\setminus B$. By [3, Lemma 4.1], there are infinitely many $v\in\mathcal{M}_{K}\setminus B$ such that there exists $n\in\mathbb{Z}_{>0}$ with $f_{v}^{n}(\tilde{o})=\tilde{o}$. For such a place $v$, we have $\tilde{o}\in P_{v}\setminus X_{v}$, which proves the lemma. ∎
@@
There exist a sequence $(x_{i})_{i=1}^{\infty}$ in ${\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and a sequence $(v_{i})_{i=1}^{\infty}$ in $\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{i}}(\mathbf{n}(\lambda(x_{i})))>0$ for all $i\geq 1$ and $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))=0$ for $j\neq i$. In particular, $((\mathbf{n}(\lambda(x_{i})))_{i=1}^{\infty},(\phi_{v_{i}})_{i=1}^{\infty})$ is an upper triangular $O$-system for $\mathbb{D}(K)_{\mathbb{Q}}$.
@@
We construct these sequences recursively.
By Lemma 6.1, $S$ is infinite. Pick $v_{1}\in S$, then there exists
$x_{1}\in Q_{v_{1}}\setminus X\subseteq{\rm Per}^{*}(f)(\overline{\mathbb{Q}}).$
We have $\phi_{v_{1}}(\mathbf{n}(\lambda(x_{1})))>0$.
Assume that we have constructed $x_{1},\dots,x_{m}\in{\rm Per}^{*}(f)(\overline{\mathbb{Q}})$ and $v_{1},\dots,v_{m}\in\mathcal{M}_{K}\setminus B$ such that $\phi_{v_{j}}(\mathbf{n}(\lambda(x_{i})))\geq 0$ and equality holds if and only if $j\neq i$. The set $\cup_{i=1}^{m}Q_{v_{i}}\setminus X$ is finite. Then there exists a finite set $T_{m}\subseteq\mathcal{M}_{K}$ such that for all $x\in\cup_{i=1}^{m}Q_{v_{i}}\setminus X$ and $v\in\mathcal{M}_{K}\setminus T_{m}$, we have $\phi_{v}(\mathbf{n}(\lambda(x)))=0$. By Lemma 6.1, we can take a place $v_{m+1}\in S\setminus(\{v_{1},\dots,v_{m}\}\cup T_{m})$. Then $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{i})))=0$ for all $1\leq i\leq m$. Pick $x_{m+1}\in Q_{v_{m+1}}\setminus X$. We have $\phi_{v_{m+1}}(\mathbf{n}(\lambda(x_{m+1})))>0$. It follows that $x_{m+1}\notin\cup_{i=1}^{m}Q_{v_{i}}$. Thus $\phi_{v_{i}}(\mathbf{n}(\lambda(x_{m+1})))=0$ for all $1\leq i\leq m$. We conclude the proof of Lemma 6.2. ∎
The conclusion now follows from Corollary 5.5 and Remark 5.6.∎
@@
Denote by $\mathcal{C}_{f}$ the set of critical points of $f$. Since $f$ is not PCF, there exists $o\in\mathcal{C}_{f}$ which is not $f$-preperiodic. Fix such a critical point $o$. Fix a subfield $K$ of $\mathbf{k}$ such that $K/\mathbb{Q}$ is finitely generated, $\tau(K)=K$, and $o,f$ are defined over $K$. Without loss of generality, we may assume that $\mathbf{k}=\overline{K}$.
Take a finitely generated $\mathbb{Z}$-subalgebra $A$ of $K$ with ${\rm Frac}(A)=K$ and $\tau(A)=A$. After shrinking ${\rm Spec}(A)$, we may assume that there exists an endomorphism $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ over $A$ whose restriction $f_{K}:\mathbb{P}^{1}_{K}\to\mathbb{P}^{1}_{K}$ on the generic fiber $\mathbb{P}^{1}_{K}$ satisfies $f=f_{K}\otimes_{K}\mathbf{k}$.
For every $c\in{\rm Spec}(A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}})(\overline{\mathbb{Q}})$, denote by $f_{c}$ the specialization of $f_{A}$ at $c$, and $o_{c}$ the specialization of $o$ at $c$. Then $o_{c}$ is a critical point of $f_{c}$. By [9, Lemma 3.3], we can take $c\in{\rm Spec}(A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}})(\overline{\mathbb{Q}})$ such that the $f_{c}$-orbit of $o_{c}$ is infinite, i.e., $o_{c}$ is not $f_{c}$-preperiodic. In particular, the specialization map $f_{c}$ is not PCF. Take a number field $L\subseteq\overline{K}$ such that $c$ is defined over $L$ and $\tau(L)=L$. Denote by $A_{1}$ the $\mathbb{Z}$-algebra generated by $A$, $\mathcal{O}_{L}$, and $\tau(\mathcal{O}_{L})$ in $A\otimes_{\mathbb{Z}}\overline{\mathbb{Q}}$. We may replace ${\rm Spec}(A)$ by a suitable non-empty affine Zariski open subset of ${\rm Spec}(A_{1})$ such that $f_{A}:\mathbb{P}^{1}_{A}\to\mathbb{P}^{1}_{A}$ remains everywhere well-defined and $c\in{\rm Spec}(A\otimes_{\mathcal{O}_{L}}L)(L)$ after the replacement. We view ${\rm Spec}(A)$ as an $\mathcal{O}_{L}$-scheme. After shrinking ${\rm Spec}(A)$, the Zariski closure of $c$ in ${\rm Spec}(A)$ is isomorphic to ${\rm Spec}(\mathcal{O}_{L,S})$ for a finite set $S\subseteq\mathcal{M}_{L}$ of places of $L$ containing all archimedean places, which corresponds to a prime ideal $p$ of $A$.